### v1

# Worldline EFT treatment of quadratic and cubic gravity theories

Abstract
This paper explores modifications to General Relativity (GR) by considering higher-order curvature terms in the gravitational action, specifically focusing on the quadratic Ricci scalar and a particular cubic contraction of the Riemann tensor. These modifications introduce new interactions at short distances, potentially altering the dynamics of compact objects. We calculate the effective two-body binding potential energy for these modified theories to quantify these effects using the worldline effective field theory (WEFT) formalism. This approach allows us to systematically integrate out short-distance gravitational effects, capturing the modifications to the binding potential. Our results demonstrate how the quadratic Ricci scalar and cubic Riemann tensor terms contribute to the two-body interaction at the leading order, highlighting deviations from classical GR predictions. These findings provide insight into the possible observational signatures of such modified gravity theories in binary systems and other astrophysical settings.
I Introduction and Motivation
Ever since [1] developed a novel method for describing the dynamics of General Relativity (GR) using the Worldline Effective Field Theory (WEFT), interest in it has spiked. Using the path integral approach of Quantum Field Theory (QFT), one can now study both conservative and radiative dynamics of the binary problem in gravity. While the calculations of the binary problem have progressed quite a lot in the standard post-Newtonian (PN) formalism, it is slow and cumbersome. WEFT of GR has delivered results for the conservative and radiative dynamics in a few years, compared to what it took for the PN formalism to achieve in some decades. This efficiency has attracted a significant amount of attention to the field of WEFTs in recent years. Many reviews of the classical aspects of WEFTs have come up, which include [2, 3, 4, 5], as well as quantum gravity effects like [6, 7, 8, 9, 10, 11].
Since the WEFT approach to GR has been tackled to a very high PN order, the next logical step is modifying the gravitational action. Modified theories of gravity have gained a lot of traction since the very early days of GR, whenever it was evident that GR was not a complete theory of gravity. The singularity at the centre of a black hole (BH), the flattening of the rotation curves of galaxies, the cosmological constant problem, etc., have plagued GR. Two approaches to dealing with the problems of GR have emerged: string theory and loop quantum gravity. At the classical limit, both theories are expected to reproduce GR. This study will deal with the first relevant subdominant classical effect of a possible high-energy theory of gravity.
GR is obtained from the Einstein-Hilbert (EH) action defined as follows
$S_{EH}=\frac{c^{4}}{16\pi G}\int d^{4}x\sqrt{-g}R$ (1)
where $R$ is the Ricci Scalar. In the renormalization program of GR, it was found by [12] that GR was renormalizable at one loop order. It was found that the divergences could be encoded in curvature squared terms like $R^{2},R_{\mu\nu}R^{\mu\nu},R_{\mu\alpha\rho\beta}R^{\mu\alpha\rho\beta}$. However, it was also found in [12] and by later authors that the curvature squared term could be field redefined away by a suitable metric redefinition. Hence, it is not prudent to include curvature squared terms for a proper theory of modified gravity in the absence of matter or BH spacetime. However, as a proof of concept, we will show the leading order change in the two-particle effective potential for the Ricci squared term added to the EH action.
It has been known for quite some time that the S matrix of GR is divergent (non-renormalizable) at two-loop order in four dimensions [13, 14]. The two loop divergences correspond to curvature cubed terms (scalars) and quadratic scalars of the covariant derivative of Ricci scalars and tensors. Utilizing the field equations makes most of the terms vanish except two Riemann cubed terms, of which only one is independent. Therefore, all two-loop divergences of the EH action can be parameterized by the following term
$\Gamma_{\infty}^{(2)}=\sqrt{-g}\,R^{\alpha\beta}_{\quad\gamma\delta}R^{\gamma\delta}_{\quad\mu\nu}R^{\mu\nu}_{\quad\alpha\beta}.$ (2)
While the particular curvature cubed term of Eq. (2) appears in the renormalization program as an actual divergence that cannot be renormalized, at a more recent time, a different procedure of field redefinition led to the same result. In [15], it was shown that field redefinitions of the metric tensor could get rid of all quadratic terms as well all but two Riemann curvature cubed terms from an action containing higher dimensional terms added to the EH action as an EFT. It was further shown in the similar spirit of [13, 14] by [16] that out of the two Riemann cubed terms, only one is independent and is again given by
$\sqrt{-g}\,\lambda_{ev}\,R^{\alpha\beta}_{\quad\gamma\delta}R^{\gamma\delta}_{\quad\mu\nu}R^{\mu\nu}_{\quad\alpha\beta}.$ (3)
where $\lambda_{ev}$ is some constant parametrizing the strength of the term in the action. This Riemann cubed term of Eq. (3) cannot be field redefined away. Thus, two seemingly different approaches seem to lead to the same Riemann cubed term, implying a connection between the renormalization program and field redefinition strategies. Hence, when added to the EH action, the Riemann cubed term can be a candidate for the first subdominant effect of a low energy limit of a high energy theory of gravity in vacuum. Therefore, in this study, we find the two-particle effective potential, using the WEFT formalism, of the EH action plus the Riemann cubed term, that is
$S=\frac{c^{4}}{16\pi G}\int d^{4}x\sqrt{-g}\left(R+\alpha\,R^{\alpha\beta}_{\quad\gamma\delta}R^{\gamma\delta}_{\quad\mu\nu}R^{\mu\nu}_{\quad\alpha\beta}\right)$ (4)
where $\alpha$ here is a fundamental constant that quantifies the cubic term’s strength. Non-rotating black hole solutions for this theory were found in [15]. The slowly rotating solution (along with the cubic odd parity part) was found in [17], and perturbations of the rotating solution were studied in [18]. The Quasi-normal mode frequencies for the non-rotating case for this particular theory and the odd parity part have been recently found in [19]. Recently, Ref. [20] studied the degrees of freedom in this model and found that the theory only carries a spin-2 field but features a preferred frame at short distances, suggesting a possible connection to unknown ultraviolet physics.
This paper is structured as follows: In Sec. II, we briefly review the WEFT formalism and reproduce the effective binding potential energy for a BBH system until the second post-Newtonian (PN) order in GR. In Sec. III, we derive the leading order change in the two-particle effective potential binding energy due to adding the Ricci squared term to the EH action. We obtain the same for the Riemann cubed gravity in Sec. IV. Finally, in Sec. V, we conclude this study by identifying future directions and discuss the implications of the results of the current study to the two body problem of gravity.
We have worked with $\hbar=1$ units throughout the text. We use the mostly plus signature of the Minkowski metric.
II Worldline EFT of GR: a mini review
Worldline effective field theories (EFTs) of gravity are a powerful framework used to study the dynamics of massive objects, such as black holes or neutron stars, in GR and modified theories of gravity. They combine quantum field theory and general relativity techniques to simplify the complex gravitational interactions between massive bodies. This particular framework has proven itself to be very efficient in obtaining various PN results of GR compared to the standard PN formalism, which required decades of research. The word ‘effective’ in EFTs of gravity or other QFTs implies that an approximation is being made when it comes to the scales (which can be energy, for example) of the problem. In theories which have phenomena covering a wide range of energy or length scales and where the various scales are decoupled from each other, EFTs are used to calculate the effects of a different energy (length) scale, usually higher, on the energy (length) scale in question, usually lower. In the case of gravity, short (long) distance effects are synonymous with high (low) energies.
The compact binary problem of gravity is one such problem where there is a clear separation of scales at the early in-spiral, or when the factor $\frac{v}{c}$ ($v$ being the characteristic orbital velocity) is close to zero (non-relativistic). In such a situation, the hierarchy of scales become
$d<<r<<\lambda$ (5)
where $d$ is the length scale of the compact objects in question, $r$ is the relative separation between them, and $\lambda\sim\frac{r}{v}$ is the wavelength of GW radiation due to the time-varying quadrupolar nature of the binary system. It is to be noted that the above scales become comparable to each other at the late inspiral stage, and the whole formalism (similar to PN) breaks down.
In the hierarchy of scales, we first encounter the physics in the region of one of the compact objects we are interested in. A compact object, at the first approximation, can be thought of as a point particle whose centre of mass is moving along a worldline $x^{\mu}(\lambda)$ that is coupled to the gravitational field $g_{\mu\nu}$, where $\lambda$ is an arbitrary affine parameter acting as some time, or can be taken as the proper time along its trajectory. However, real compact objects in nature are not point particles but have some ‘spread’. This spread can be perturbatively treated as multipole moments of the source energy-momentum tensor. This requires an in-depth knowledge of the source energy-momentum tensor. However, we generally do not have complete information on the energy-momentum tensor. In such a case, we can parametrize our ignorance about the spread of the compact object as a series of operators with Wilson coefficients proportional to the so-called ‘Love numbers’ of the compact object. This spread and the Love numbers are directly related to the tidal response of one of the compact objects in the presence of the other. Thus, the effective point particle action can be written as follows
[…]
where in the first term $d\tau=\sqrt{g_{\mu\nu}dx^{\mu}dx^{\nu}}$, and is the result of the geodesic motion with respect to the metric $g_{\mu\nu}$. The second and the third term with Wilson coefficients $c_{S/T}$ can be removed by a field redefinition of the metric, as was argued in [1]. Therefore, they can be removed. The third and fourth terms contain contractions of the so-called electric and magnetic parts of the Riemann tensor defined as follows:
$E_{\mu\nu}=R_{\mu\nu\alpha\beta}\dot{x}^{\alpha}\dot{x}^{\beta}\qquad B_{\mu\nu}=\epsilon_{\mu\alpha\beta\rho}R^{\alpha\beta}_{\sigma\nu}\dot{x}^{\sigma}$ (7)
However, since we are only interested in BBH space-times in this study, the constants $c_{E}$ and $c_{B}$ also vanish, as was shown in [21]. Hence, we are simply left with the first term, that is
[…]
where $x^{\mu}_{pp}(\Lambda)$ is the worldline of one of the point particles parametrized by $\Lambda$ and $v^{\mu}=\frac{dx^{\mu}}{d\Lambda}$.
A step up from the smallest length scale $d$ is the near zone $r$. In this zone, the weak field regime applies, and one can expand the metric tensor about the flat spacetime $\eta_{\mu\nu}$ as
$g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$ (9)
Therefore, to obtain the effective two-particle PN action (and hence the Lagrangian or the effective binding potential), one needs to integrate out the graviton $h_{\mu\nu}$ in the following manner
$e^{iS_{eff}}=\int Dh_{\mu\nu}e^{iS_{EH}[h]+iS_{GF}[h]+iS_{pp}[h,x^{\mu}]}.$ (10)
where
$S_{EH}=\frac{c^{4}}{16\pi G}\int d^{4}x\sqrt{-g}R[h,\partial h,\partial^{2}h]$ (11)
is the Einstein-Hilbert action and
$S_{GF}=-\frac{c^{4}}{32\pi G}\int d^{4}x\sqrt{-g}\Gamma_{\mu}\Gamma^{\mu}$ (12)
where $\Gamma^{\mu}=\Gamma^{\mu}_{\alpha\beta}g^{\alpha\beta}$ is the contracted Christoffel symbol and $S_{GF}$ is known as a gauge-fixing term. We use the harmonic gauge throughout this study.
To fully exploit the separation of scales in the binary problem, one usually splits further the perturbation $h_{\mu\nu}$ into potential $(H_{\mu\nu})$, corresponding to the short length scale, and radiation $(\bar{h}_{\mu\nu})$ modes, corresponding to the larger length scale, as was done in [3] and [5]. This separates the physics of $r$ from the physics of $\lambda$. The potential modes are treated as near instantaneous, given that the rate of change of the configuration of the binary is much slower than the rate at which gravitons are exchanged between the two BHs. To get the conservative dynamics, or the binding potential of the binary, one needs to integrate out both the potential and radiation mode, that is
$e^{iS_{eff}[x^{\mu}]}=\int D\bar{h}_{\mu\nu}\int DH_{\mu\nu}e^{iS_{tot}(x^{\mu},\bar{h},H)},$ (13)
where $S_{tot}$ was defined in the RHS of Eq. (10).
Given the nonlinearity of the net action in the RHS of Eq. (10) and (13), one must perform the integral perturbatively using Feynman diagrams. Once the effective action has been calculated, the real part of the action will lead to the conservative Lagrangian of the two-point particle system with the graviton integrated out. Similarly, the imaginary part of the effective action encodes the system’s dissipative part or is proportional to the total radiated power through GW emission. However, in this paper, we will restrict ourselves to the conservative part of the action only and deal with the radiative effects in a follow-up paper.
This paper, however, does not use the decomposition of the potential+radiation method of the metric perturbation; instead, it uses a Kaluza-Klein decomposition of the metric/metric perturbation. Given a d+1 dimensional spacetime, we can rewrite the metric components, as was done in [22], as follows
$g_{\mu\nu}=e^{2\phi}\Biggl{(}\begin{matrix}-1&A_{j}\\ A_{i}&e^{-c_{d}\phi}\gamma_{ij}-A_{i}A_{j}\end{matrix}\Biggr{)}\ ,\qquad\qquad\gamma_{ij}=\left(\delta_{ij}+\sigma_{ij}\right)\ .$ (14)
where $c_{d}=2\left(\frac{d-1}{d-2}\right)$ and $\delta_{ij}$ is the flat d dimensional metric. Using the full metric in Eq. (14) and the metric perturbation equation (9), we can read off the metric perturbation $h_{\mu\nu}$ as
$h_{\mu\nu}=\Biggl{(}\begin{matrix}1-e^{2\phi}&A_{j}\\ A_{i}&e^{(2-c_{d})\phi}\gamma_{ij}-e^{2\phi}A_{i}A_{j}-\delta_{ij}\end{matrix}\Biggr{)}$ (15)
where $\phi$ is a scalar field, $A^{i}$ is a d-dimensional vector field, and $\sigma_{ij}$ is a $d\times d$ symmetric tensor field. Given the three (scalar, vector, and tensor) fields, Eq. (10) or (13) gets modified into the following
$e^{iS_{eff}(x^{\mu})}=\int D\phi DA_{i}D\sigma_{jk}e^{iS_{EH}[\phi,A_{i},\sigma_{jk}]+iS_{pp}[x^{\mu},\phi,A_{i},\sigma_{jk}]+iS_{GF}[\phi,A_{i},\sigma_{jk}]}$ (16)
We obtain the interaction terms for the perturbative path integral treatment by expanding the three terms of the action $(S_{EH},S_{pp},S_{GF})$ till the fifth order in $h_{\mu\nu}$ and then decompose $h$ into $\left(\phi,A_{i},\sigma_{jk}\right)$, as shown in Eq. (15). We also scale the fields $\left(\phi,A_{i},\sigma_{jk}\right)$ in powers of $c$ to obtain a PN series of the interaction terms in the bulk and worldline action in the following manner
$\phi$$\rightarrow$ $\frac{\phi}{c^{2}}$ (17)
$A_{i}$$\rightarrow$ $\frac{A_{i}}{c^{3}}$ (18)
$\sigma_{jk}$$\rightarrow$ $\frac{\sigma_{jk}}{c^{4}}.$ (19)
After this, we perturbatively evaluate Eq. (16) by first identifying the appropriate two-point correlation functions (or the propagators) and the interaction vertices of the fields $\left(\phi,A_{i},\sigma_{jk}\right)$ from the expansion of the full action in Kaluza-Klein (Kol-Smolkin) variables and by transforming to the Fourier domain. The path-integral is computed diagrammatically, order by order, characterized by a definite scaling of the speed of light $c$ and Newton’s constant $G$. Since we first look for the non-relativistic limit (or slow motion of the sources), we can take $\frac{k_{0}}{|\vec{k}|}$ as small and expand the propagators around $\frac{k_{0}}{|\vec{k}|}=0$. For the exact form of the full propagators, see Appendix A of Ref. [5].
[…]
We have found that the 2 PN two-particle effective Lagrangian does not change after including the fifth order of perturbation of the metric in the action, as well as reproduces the same from previous literature [23], where they quote effective Lagrangian for only one of the particles. We reproduce the two-particle effective binding potential energy $V_{eff}$ (after summing over both the particles) till 2 PN in this paper for the benefit of the reader as follows:
$V_{eff}$ $=$ $V_{N}+\frac{1}{c^{2}}V_{1PN}+\frac{1}{c^{4}}V_{2PN}$ (20)
$V_{N}$ $=$ $-\frac{Gm_{1}m_{2}}{r}$ (21)
$V_{1PN}$ $=$ $-Gm_{1}m_{2}\left[\frac{\left\{3v_{1}^{i}v_{1,i}+\left(3v_{2}^{i}-7v_{1}^{i}\right)v_{2,i}\right\}}{2r}-\frac{v_{1}^{i}x_{i}v_{2}^{j}x_{j}}{2r^{3}}\right]+\frac{G^{2}m_{1}m_{2}\left(m_{1}+m_{2}\right)}{2r^{2}}$
$V_{2PN}$ $=$ $-Gm_{1}m_{2}\left[-\frac{1}{8r^{3}}\left\{v_{1}^{i}\left[v_{1}^{j}x_{j}\left\{x_{k}\left(6v_{2}^{k}v_{1,i}-12v_{2}^{k}v_{2,i}-a_{2}^{k}x_{i}\right)+x_{i}v_{2}^{k}v_{2,k}\right\}+v_{2}^{j}v_{2}^{k}x_{k}\left(6x_{i}v_{2,j}+x_{j}v_{1,i}\right)\right]\right.\right.$ (23)
$\left.+a_{1}^{i}x_{i}v_{2}^{j}x_{j}v_{2}^{k}x_{k}\right\}+\frac{1}{8r}\left\{a_{1}^{i}x_{i}v_{2}^{j}v_{2,j}-14a_{1}^{i}v_{2}^{j}x_{j}v_{2,i}+v_{1}^{i}v_{1,i}\left(7v_{1}^{j}v_{1,j}-10v_{1}^{j}v_{2,j}+3v_{2}^{j}v_{2,j}\right.\right.$
$\left.-a_{2}^{j}x_{j}\right)+2v_{1}^{i}\left[v_{2}^{j}\left(-6x_{i}a_{2,j}+6x_{j}a_{1,i}-5v_{2,i}v_{2,j}\right)+v_{1}^{j}\left(7x_{j}a_{2,i}+v_{2,i}v_{2,j}\right)\right]+7v_{2}^{i}v_{2}^{j}v_{2,i}v_{2,j}$
$\left.\left.-a_{1}^{i}a_{2}^{j}x_{i}x_{j}\right\}+\frac{15}{8}ra_{1}^{i}a_{2,i}+\frac{3v_{1}^{i}x_{i}v_{1}^{j}x_{j}v_{2}^{k}x_{k}v_{2}^{l}x_{l}}{8r^{5}}\right]-G^{2}m_{1}m_{2}\left[\frac{1}{4r^{2}}\left\{m_{1}\left[8v_{1}^{i}v_{1,i}+7\left(v_{2}^{i}-2v_{1}^{i}\right)v_{2,i}\right]\right.\right.$
$\left.\left.+m_{2}\left(7v_{1}^{i}v_{1,i}+2\left(4v_{2}^{i}-7v_{1}^{i}\right)v_{2,i}\right)\right\}+\frac{1}{2r^{4}}\left\{x_{i}x_{j}\left(m_{1}v_{2}^{i}v_{2}^{j}+m_{2}v_{1}^{i}v_{1}^{j}\right)\right\}\right]$
$-\frac{G^{3}m_{1}m_{2}}{2r^{3}}\left(m_{1}^{2}+6m_{2}m_{1}+m_{2}^{2}\right)$
where $r=\sqrt{x^{i}x_{i}}$ is the distance between the two point particles, $x^{i}$ is the separation vector, $a^{i}_{K}$ for $K=1,2$ is the acceleration vector and $v^{i}_{K}$ for $K=1,2$ is the velocity vector for the first and the second object. We only give the effective potential and not the PN kinetic energy since it does not change under the change in the bulk action. In the rest of this study, we keep the worldline unchanged while we change the bulk action.
III Proof of concept: quadratic gravity
We start by adding the Ricci scalar squared to the EH action and by fixing the gauge to harmonic. The full action in the harmonic gauge is then given in d = 3 as follows
$S_{quad}$ $=$ $S_{bulk}+S_{pp}$ (24)
$S_{bulk}$ $=$ $\frac{c^{4}}{16\pi G}\int d^{4}x\,\sqrt{-g}\,(R+\alpha R^{2}-\frac{1}{2}\Gamma^{\mu}\Gamma_{\mu})$ (25)
[…]
where $\Gamma^{\mu}$ was defined right after Eq. (12), $m_{a}$ for $a=1,2$ are the rest masses of two point particles moving on a wordline $x^{\mu}_{a}$.
While $S_{pp}$ remains the same as in GR, $S_{bulk}$ changes due to the inclusion of the $R^{2}$ term. The changes in the bulk action on perturbing the same and decomposing into Kaluza-Klein variables, as given in Eq. (15), are given till 1 PN as follows (we include the GR terms as well):
$S_{bulk}$ $=$ $\frac{1}{G}\int d^{4}x\hskip 2.84544pt\left[\frac{\phi\nabla^{2}\phi}{8\pi}-\alpha\frac{\nabla^{2}\phi\nabla^{2}\phi}{4\pi}+\frac{1}{c^{2}}\left\{\frac{4(\partial_{t}\phi)^{2}-A_{i}\nabla^{2}A^{i}}{32\pi}+\frac{\alpha}{4\pi}\left(2\nabla^{2}\phi(\partial\phi)^{2}-4\partial^{i}\phi\partial_{i}\partial_{j}\phi\partial^{j}\phi\right.\right.\right.$ (27)
$\left.\left.\left.+2\phi\partial_{i}\nabla^{2}\phi\partial^{i}\phi+6\nabla^{2}\phi(\partial_{t}\phi)^{2}-2\partial^{i}\nabla^{2}\phi\partial_{t}A_{i}-\partial_{i}\nabla^{2}\sigma\partial^{i}\phi-\nabla^{2}\phi\partial_{k}\partial_{j}\sigma^{jk}\right)\right\}+\mathcal{O}\left(\frac{1}{c^{4}}\right)\right]$
where $\sigma=\sigma^{i}_{i}$, $\nabla^{2}\equiv\delta^{ij}\partial_{i}\partial_{j}$, $\delta_{ij}$ is the 3 dimensional Minkowski spatial metric, and $(\partial\phi)^{2}=\partial_{i}\phi\partial^{i}\phi$. We also give the Kaluza-Klein decomposed point particle worldline action for one of the particles till 2 PN for the reader’s convenience
$S_{pp}$ $=$ $\int dt\left\{-mc^{2}+\frac{1}{2}mv^{2}-m\phi+\frac{m}{c^{2}}\left(\vec{A}\cdot\vec{v}+\frac{1}{8}|\vec{v}|^{4}-\frac{3}{2}|\vec{v}|^{2}\phi-\frac{1}{2}\phi^{2}\right)\right.$
$\left.+\frac{m}{c^{4}}\left[\frac{1}{2}|\vec{v}|^{2}\vec{A}\cdot\vec{v}+\frac{1}{16}|\vec{v}|^{6}+\frac{1}{2}v^{i}v^{j}\sigma_{ij}+\left(\vec{A}\cdot\vec{v}-\frac{7}{8}|\vec{v}|^{4}\right)\phi+\frac{9}{4}|\vec{v}|^{2}\phi^{2}-\frac{1}{6}\phi^{3}\right]+\mathcal{O}\left(\frac{1}{c^{5}}\right)\right\}$
III.1 Newtonian order or 0 PN
At the Newtonian order, because of the particular scaling of Eq. (17), we only get $\phi$ and derivatives in Eq. (27). Taking the Fourier transform of $\phi$ one can write $\phi(x)$ in terms of $\phi(k)$ as
$\phi(x^{\mu})$ $=$ $\int\frac{d^{4}k}{(2\pi)^{4}}e^{ik_{\mu}x^{\mu}}\phi(k^{\mu})=\int\frac{d^{4}k}{(2\pi)^{4}}e^{i(\vec{k}\cdot\vec{x}-ck_{0}t)}\phi(k^{\mu})$ (29)
In the non-relativistic limit, at the leading order, $k_{0}$ is ignored. The 0 PN diagram is just the exchange of one graviton (or the scalar part of the Kaluza-Klein metric decomposition) between the two worldlines and is given as follows.
In the above diagram, a double line corresponds to a worldline factor, whereas the thin line corresponds to the $\phi$ propagator. The above diagram corresponds to the leading order term in the series expanded generating functional $Z[J]$. The generating functional in quantum field theory encapsulates all correlation functions of a field theory and is defined through a path integral formulation, enabling the systematic computation of n-point functions via functional derivatives with respect to a source term [24]. $Z[J]$ is defined as follows (since the diagram only involves $\phi$ we have suppressed the $A$ and $\sigma$ fields in the following)
$Z[J]=\int D\phi e^{iS[\phi(x^{\mu})]+J(x^{\mu})\phi(x^{\mu})}$ (30)
where at the leading order,
$S[\phi(x^{\mu})]\approx\frac{1}{8\pi G}\int d^{4}x\left(\phi\nabla^{2}\phi-\alpha\nabla^{2}\phi\nabla^{2}\phi\right)$ (31)
and $J(x^{\mu})$ is the leading order source/interaction term coming from $S_{pp}$ and is given as follows
[…]
where $x^{\mu}_{a}$ is the worldline of the $a$-th particle. As was performed in Ref. [25], the ‘free’ part of the action is kept in the exponential, while the source/interaction term (or the $J$ term) is expanded in a series. In the Fourier domain, the series expanded generating functional is given as follows
$Z[\tilde{J}]$$\approx$ $\int D\phi(k)\,\,e^{{-\frac{i}{8\pi G}\int d^{4}k\phi(-k)(|\vec{k}^{2}|+2\alpha|\vec{k}|^{4})\phi(k)}}\left(\frac{1}{2}\int d^{4}k_{1}d^{4}k_{2}\tilde{J}(-k_{1})\tilde{J}(-k_{2})\phi(k_{1})\phi(k_{2})\right)$
where
[…]
It is to be noted that although there is a double sum in Eq. (III.1) owing to the product of the two $J$’s, we only take cross terms into account while performing the calculations and hence, we ignore the self energy terms.
The propagator is obtained at the leading order, and it is to be noted that it changes due to the term proportional to $\alpha$, and hence it changes the 2-point function as well
$<\phi(k_{1})\phi(k_{2})>=\frac{-4\pi iG(2\pi)^{4}\delta^{4}(k_{1}+k_{2})}{(k_{1}^{2}+2\alpha k_{1}^{4})}$ (35)
We then calculate the amplitude for the one graviton exchange and take its inverse Fourier transform to obtain the effective potential, that is
$V_{eff}$ $\propto$ $-\int d^{4}k_{1}d^{4}k_{2}\tilde{J}(-k_{1})\tilde{J}(-k_{2})<\phi(k_{1})\phi(k_{2})>$ (36)
$=$ $-\frac{Gm_{1}m_{2}}{r}\left(1-e^{\frac{-r}{\sqrt{2\alpha}}}\right)$ (37)
An important integration that we use in the calculation is
$\int\frac{d^{3}k}{(2\pi)^{3}}\frac{e^{i\vec{k}\cdot\vec{r}_{a}}}{|\vec{k}|^{2}+2\alpha|\vec{k}|^{4}}=\frac{1-e^{\frac{-r}{\sqrt{2\alpha}}}}{4\pi r}$ (38)
The result of Eq. (36) exactly matches what was found in Section 8 of [10]. We can also consider $\alpha$ as a negative parameter; in this case, the above diagram gives
$V_{eff}=-\frac{Gm_{1}m_{2}}{r}\left[1-\cos{\left(\frac{r}{\sqrt{2\alpha}}\right)}\right]$ (39)
III.2 PN1 : Alpha as a perturbing parameter
If we take $\alpha$ as a perturbing parameter, there is no change at 0 PN since the Yukawa potential becomes the Newtonian potential plus a delta function. The delta function will only contribute when the two particles are at the same position, which is not the situation we are looking at. Hence, we can ignore the delta function. Now the interaction terms that we have to consider are at 1 PN and are given by Eq. (27), which leads to the following diagrams with non-zero contribution to the end result
The series expanded generating functional now becomes
$Z$$\approx$ $\frac{1}{2}\int D\phi(k)e^{{-\frac{i}{8\pi G}\int d^{4}k\hskip 2.84544pt\phi(-k)\,\,k^{2}\,\,\phi(k)}}$
$\times$ $\int_{k_{1}...k_{6}}J(-k_{4})J(-k_{5})J(-k_{6})\phi(k_{1})\phi(k_{2})\phi(k_{3})\phi(k_{4})\phi(k_{5})\phi(k_{6})\vec{k_{2}}\cdot\vec{k_{3}}\leavevmode\nobreak\ k_{1}^{2}\delta(\vec{k_{1}}+\vec{k_{2}}+\vec{k_{3}})$
which implies that the amplitude of the $\phi^{3}$ interactions (as in the above diagrams) become
$\mathcal{M}$ $\propto$ $\frac{\alpha}{2}\int_{k_{1}...k_{6}}J(-k_{4})J(-k_{5})J(-k_{6})\left<\phi(k_{1})\phi(k_{2})\phi(k_{3})\phi(k_{4})\phi(k_{5})\phi(k_{6})\right>\vec{k_{2}}\cdot\vec{k_{3}}\leavevmode\nobreak\ k_{1}^{2}\delta(\vec{k_{1}}+\vec{k_{2}}+\vec{k_{3}})$
where $\int_{k_{1}...k_{6}}\equiv\int\prod_{i=1}^{6}\frac{d^{4}k_{i}}{(2\pi)^{4}}$. Performing the integral and using the combinatoric factors, the change in the PN1 result is given as follows
$V_{eff,\alpha}$ $=$ $-\frac{2\alpha G^{2}}{c^{2}}\frac{m_{1}m_{2}(m_{1}+m_{2})}{r_{12}^{4}},$ (42)
where $r_{12}$ is the inter orbital separation. For negative $\alpha$, the result changes by an overall sign.
IV The two-particle effective Lagrangian of cubic gravity
In this section, we try to calculate the leading order change in the two-body effective potential by integrating the gravitational degrees of freedom from the following action
$S_{cubic}$ $=$ $S_{bulk}+S_{pp}$ (43)
$S_{bulk}$ $=$ $\frac{c^{4}}{16\pi G}\int d^{4}x\left(R+\alpha\,R^{\alpha\beta}_{\quad\gamma\delta}R^{\gamma\delta}_{\quad\mu\nu}R^{\mu\nu}_{\quad\alpha\beta}-\frac{1}{2}\Gamma^{\mu}\Gamma_{\mu}\right)$ (44)
where $S_{pp}$ remains the same as in Eq. (26). Like in the previous section, we use the harmonic gauge fixing term. Upon perturbing the metric and decomposing the metric perturbation as in Eq. (15), we find the following interaction terms in two different PN orders:
IV.1 PN-1
At the first PN order, the bulk interaction terms are given as follows:
$\alpha\left[\frac{3\left(\partial_{i}\partial_{i}\phi\right)\left(\partial_{j}\partial_{l}\phi\partial_{j}\partial_{l}\phi\right)}{2G\pi}-\frac{\left(\partial_{i}\partial_{j}\phi\,\partial_{j}\partial_{l}\phi\,\partial_{l}\partial_{i}\phi\right)}{G\pi}\right],$ (45)
which correspond to the following diagrams
Inverse Fourier transforms of the amplitudes of the above two diagrams either vanish or become delta functions with the inter-orbital separation as an argument. Hence, there is no change in the two-body effective potential at the first PN order. We move on to the second PN order.
IV.2 PN-2
In the second PN order, we look at contributions from different Klauza-Klein fields appearing in the bulk interaction. We give the following: the type and kind of fields (or their mixtures) appearing as interaction terms, the interaction terms themselves, and the respective Feynman diagrams when coupled to the worldlines. In the following, double lines denote worldline factors, thin lines denote the $\phi$ propagator, dashed lines denote the $A^{i}$ propagator, and finally, wavy lines denote the $\sigma_{ij}$ propagator. A cross in the middle of a thin line implies that a factor of $\frac{k_{0}}{|\vec{k}|}$ has been inserted in the $\phi$ propagator.
$\underline{\phi^{4}}$
$\frac{-3(\partial_{i}\partial^{i}\phi)^{2}\leavevmode\nobreak\ (\partial_{j}\phi\partial^{j}\phi)}{2G\pi}+\frac{3(\partial_{i}\partial^{i}\phi)\leavevmode\nobreak\ (\partial^{j}\phi\partial_{k}\leavevmode\nobreak\ \partial_{j}\phi\partial^{k}\phi)}{G\pi}+\frac{3(\partial_{i}\phi\partial^{i}\phi)\leavevmode\nobreak\ (\partial_{j}\partial_{k}\phi\leavevmode\nobreak\ \partial^{j}\partial^{k}\phi])}{2G\pi}$
$-\frac{15(\partial_{i}\partial^{j}\phi\leavevmode\nobreak\ \partial^{i}\phi\partial^{k}\phi\leavevmode\nobreak\ \partial_{j}\partial_{k}\phi)}{2G\pi}+\phi\left(\frac{6\partial_{k}\partial^{k}\phi\partial_{i}\partial_{j}\phi\partial^{i}\partial^{j}\phi-4\partial_{i}\partial_{j}\phi\partial^{j}\partial_{k}\phi\partial^{i}\partial^{k}\phi}{G\pi}\right)$ (46)
$\underline{\phi^{3}}$
$-\frac{7(\partial_{i}\partial_{t}\phi\leavevmode\nobreak\ \partial_{j}\partial_{t}\phi\leavevmode\nobreak\ \partial^{i}\phi)}{2G\pi}+\frac{\partial^{i}\partial_{t}\phi\leavevmode\nobreak\ \partial_{j}\partial_{i}\phi\leavevmode\nobreak\ \partial^{j}\partial_{t}\phi}{2G\pi}-3\frac{\partial_{i}\partial_{j}\phi\partial^{i}\partial^{j}\phi\partial_{t}^{2}\phi}{2G\pi}$ (47)
$\underline{A\phi^{2}}$
$-\frac{3(\partial_{i}\partial_{t}\phi\leavevmode\nobreak\ \partial^{i}\partial^{j}\phi\leavevmode\nobreak\ \partial_{j}\partial_{k}A^{k}}{4G\pi}-\frac{3(\partial_{i}\partial_{t}A^{j}\leavevmode\nobreak\ \partial^{i}\partial^{k}\phi\leavevmode\nobreak\ \partial_{j}\partial_{k}\phi)}{4G\pi}+\frac{3(\partial_{i}\partial^{j}\phi\leavevmode\nobreak\ \partial^{k}\partial_{k}A^{i}\partial_{j}\partial_{t}\phi)}{4G\pi}-\frac{3(\partial^{j}\partial_{i}A^{i}\partial_{j}\partial^{k}\phi\leavevmode\nobreak\ \partial_{k}\partial_{t}\phi)}{4G\pi}$
$+\frac{3(\partial^{j}\partial_{j}\leavevmode\nobreak\ A^{i}\leavevmode\nobreak\ \partial_{k}\partial_{t}\phi\leavevmode\nobreak\ \partial^{k}\partial_{i}\phi)}{4G\pi}-\frac{3(\partial_{i}\partial_{t}A_{j}\partial^{k}\partial^{i}\phi\leavevmode\nobreak\ \partial^{j}\partial_{k}\phi)}{4G\pi}+\frac{(\partial_{j}\partial_{t}\phi\leavevmode\nobreak\ \partial_{k}\partial_{i}\phi\leavevmode\nobreak\ \partial^{j}\partial^{k}A^{i})}{4G\pi}+\frac{(\partial_{j}\partial_{i}\phi\leavevmode\nobreak\ \partial_{k}\partial_{t}\phi\leavevmode\nobreak\ \partial^{j}\partial^{k}A^{i}}{4G\pi}$
$-\frac{(\partial^{j}\partial^{i}\phi\leavevmode\nobreak\ \partial_{i}\partial_{k}A_{j}\leavevmode\nobreak\ \partial^{k}\partial_{t}\phi)}{4G\pi}-\frac{(\partial^{j}\partial^{i}\phi\leavevmode\nobreak\ \partial_{j}\partial_{k}A_{i}\leavevmode\nobreak\ \partial^{k}\partial_{t}\phi)}{4G\pi}$ (48)
$\underline{\sigma\phi^{2}}$
$-\frac{(\partial^{i}\partial^{j}\phi\leavevmode\nobreak\ \partial_{i}\partial_{m}\sigma^{k}_{\leavevmode\nobreak\ k}\leavevmode\nobreak\ \partial^{m}\partial_{j}\phi)}{4G\pi}+\frac{3(\partial^{i}\phi\partial^{j}\phi\leavevmode\nobreak\ \partial_{k}\partial_{j}\phi\leavevmode\nobreak\ \partial^{k}\partial_{i}\phi)}{2G\pi}-\frac{(\partial^{j}\partial^{i}\phi\leavevmode\nobreak\ \partial_{i}\partial_{m}\sigma^{k}_{\leavevmode\nobreak\ k}\leavevmode\nobreak\ \partial^{m}\partial_{j}\phi)}{2G\pi}$
$+\frac{3(\partial^{j}\partial^{i}\phi\leavevmode\nobreak\ \partial^{k}\partial_{j}\phi\leavevmode\nobreak\ \partial_{i}\partial_{m}\sigma_{k}^{\leavevmode\nobreak\ m})}{4G\pi}+\frac{3(\partial^{j}\partial^{i}\phi\leavevmode\nobreak\ \partial^{k}\partial_{j}\phi\leavevmode\nobreak\ \partial_{k}\partial_{m}\sigma_{i}^{\leavevmode\nobreak\ m})}{4G\pi}-\frac{3(\partial^{j}\partial^{i}\phi\leavevmode\nobreak\ \partial^{k}\partial_{j}\phi\leavevmode\nobreak\ \partial^{m}\partial_{m}\sigma_{ik})}{4G\pi}$
$+\frac{3(\partial^{i}\partial^{m}\phi\leavevmode\nobreak\ \partial^{k}\partial^{l}\phi\leavevmode\nobreak\ \partial_{i}\partial_{l}\sigma_{km})}{2G\pi}-\frac{3(\partial^{i}\partial^{l}\phi\leavevmode\nobreak\ \partial^{k}\partial^{m}\phi\leavevmode\nobreak\ \partial_{i}\partial_{l}\sigma_{km})}{2G\pi}$ (49)
$\underline{A^{2}\phi}$
$-\frac{3(\partial_{j}\partial^{k}\phi(\partial^{m}\partial^{j}A^{i})\partial_{i}\partial_{k}A_{m})}{8G\pi}+\frac{3(\partial^{i}\partial_{l}A_{m}(\partial^{m}\partial^{l}A^{j})\partial_{i}\partial_{j}\phi)}{8G\pi}-\frac{3((\partial_{i}\partial_{l}A^{m})\partial^{i}\partial^{l}A^{j}\partial_{m}\partial_{j}\phi)}{8G\pi}$
$+\frac{3(\partial_{i}\partial_{l}A_{m}\partial^{l}\partial^{k}A^{i}\partial^{m}\partial_{k}\phi)}{8G\pi}$ (50)
Scalar propagator expanded PN 1 $\phi^{3}$ appearing at PN 2
We obtain the final result by summing over all the previous contributions at 2 PN and taking an inverse Fourier transform, leading to a change in the two-body effective potential energy due to the Riemann cubed term in the bulk gravitational action compared to GR. At 2 PN, the change is given as follows:
$V_{eff,\alpha}=$ $-\frac{24G^{2}m_{1}m_{2}}{r_{12}^{6}}\bigg{\{}m_{1}\bigg{[}3(a_{1}\cdot r_{12})-(a_{2}\cdot r_{12})\bigg{]}-m_{2}\bigg{[}3(a_{2}\cdot r_{12})-(a_{1}\cdot r_{12})\bigg{]}\bigg{\}}$
$-\frac{36G^{2}m_{1}m_{2}(m_{1}+m_{2})}{r_{12}^{6}}\left(v_{1}^{2}+v_{2}^{2}-2v_{1}\cdot v_{2}\right)+\frac{160\,G^{3}\,m_{1}\,m_{2}\left(m_{1}^{2}+m_{2}^{2}\right)}{7r_{12}^{7}}$
$+\frac{108\,G^{2}\,m_{1}\,m_{2}\left(m_{1}+m_{2}\right)}{r_{12}^{8}}\left((v_{1}\cdot r_{12})-(v_{2}\cdot r_{12})\right)^{2}+\mathcal{O}\left(\frac{1}{r_{12}^{11}}\right)$ (51)
where $r_{12}$ is the inter orbital separation, $v_{i}$ for $i=1,2$ are the velocities of BH 1 and BH 2 respectively, and $a_{i}$ for $i=1,2$ are the accelerations of BH 1 and 2 respectively. Eq. (IV.2) can be rewritten in a form where instead of individual accelerations and velocities, we use relative values. The form is given as follows
$V_{eff,\alpha}$ $=$ $-\frac{24G^{2}\mu M}{r_{12}^{6}}\bigg{[}2\delta M\vec{a}_{CM}-4\mu\vec{a}_{rel}-M\vec{a}_{rel}\bigg{]}\cdot\vec{r}_{12}-\frac{36G^{2}\mu M^{2}}{r_{12}^{6}}v_{rel}^{2}$ (52)
$+\frac{108\,G^{2}\,\mu M^{2}}{r_{12}^{8}}\left(\vec{v}_{rel}\cdot\vec{r}_{12}\right)^{2}+\frac{160\,G^{3}\,\mu M^{2}\left(M-2\mu\right)}{7r_{12}^{7}}$
where $\frac{1}{\mu}=\frac{1}{m_{1}}+\frac{1}{m_{2}}$, $M=m_{1}+m_{2}$, $\delta=\frac{m_{1}-m_{2}}{M}$, $\vec{a}_{CM}=\frac{m_{1}\vec{a}_{1}+m_{2}\vec{a}_{2}}{M}$, $\vec{a}_{rel}=\vec{a}_{2}-\vec{a}_{1}$, and $\vec{v}_{rel}=\vec{v}_{2}-\vec{v}_{1}$.
V Conclusions and Discussion
This study explores the two-particle effective potential energy for two modified gravity theories: quadratic gravity and cubic gravity. We integrated the gravitational degrees of freedom using the Worldline Effective Field Theory (WEFT) approach, employing a Kaluza-Klein metric decomposition to simplify the interactions. The analysis showed that higher-order curvature terms—specifically the quadratic Ricci scalar and a particular cubic Riemann tensor contraction—modify the gravitational binding potential at short distances, deviating from the predictions of General Relativity (GR).
For the quadratic gravity case, we observed that the potential energy correction is proportional to the inverse of the fourth power of the separation distance between two black holes. Meanwhile, the cubic gravity term contributed a correction that scaled inversely with the fifth power of the separation. These results suggest that higher-order curvature corrections could significantly alter short-distance gravitational dynamics, potentially offering a means to test these theories in highly compact systems like binary black hole mergers.
The cubic gravity modification, particularly, is noteworthy due to its connection with the renormalization program and field redefinition strategies in quantum gravity. Both approaches independently arrive at the same Riemann cubic term, implying that this term may play a fundamental role in extending GR. Future work should explore this connection further to clarify the implications for quantum gravity theories.
V.1 Future directions
While this paper focused on the conservative effects of modified gravity theories, radiative effects also play a crucial role, especially in systems that emit gravitational waves. In a follow-up study, we will investigate the radiative corrections to the effective potential and calculate the leading-order changes in gravitational waveforms due to these higher-order curvature terms. This would be an important step in understanding how modified gravity theories might manifest in astrophysical observations, particularly in the context of gravitational wave detections from binary mergers.
VI Acknowledgments
We have partially used the package EFTofPNG [26] in our calculations. We are also indebted to Raj Patil and Jan Steinhoff for important discussions and help regarding the EFTofPNG package.
Appendix A Extra diagrams for cubic gravity at 2 PN
There were eight diagrams that the EFTofPNG package could not solve. We manually calculated them and only tried to gauge the falloff behaviour of some terms. We restricted ourselves to some terms since there was a dramatic proliferation of terms. We found the following
$\phi^{3}$ $2^{nd}$ order
$\begin{split}V_{eff,\alpha}&\propto\int_{k_{1}...k_{5}}d^{3}k_{1}...d^{3}k_{5}\left[\frac{k_{1}^{2}\leavevmode\nobreak\ (k_{2}.k_{3})^{2}\leavevmode\nobreak\ k_{3}^{2}(k_{4}.k_{5})^{2}}{k_{1}^{2}\leavevmode\nobreak\ k_{2}^{2}\leavevmode\nobreak\ k_{3}^{2}\leavevmode\nobreak\ k_{4}^{2}\leavevmode\nobreak\ k_{5}^{2}}\right]e^{ik_{1}.x_{1}}e^{-i(k_{2}+k_{3}+k_{5}).x_{2}}\delta^{3}(k_{1}-k_{2}-k_{3})\delta^{3}(k_{3}-k_{4}-k_{5})\\ &\text{integrate over $k_{1}$ and $k_{3}$}\\ &=\int_{k_{2}k_{4}k_{5}}d^{3}k_{2}\leavevmode\nobreak\ d^{3}k_{4}\leavevmode\nobreak\ d^{3}k_{5}\left[\frac{((k_{2}.(k_{4}+k_{5}))^{2}\leavevmode\nobreak\ (k_{4}.k_{5})^{2}}{k_{2}^{2}\leavevmode\nobreak\ k_{4}^{2}\leavevmode\nobreak\ k_{5}^{2}}\right]e^{i(k_{2}+k_{4}+k_{5}).r}\\ &=\int_{k_{2}k_{4}k_{5}}d^{3}k_{2}\leavevmode\nobreak\ d^{3}k_{4}\leavevmode\nobreak\ d^{3}k_{5}\left[\left(\frac{(k_{2}.k_{4})^{2}(k_{4}.k_{5}))^{2}}{k_{2}^{2}\leavevmode\nobreak\ k_{4}^{2}\leavevmode\nobreak\ k_{5}^{2}}\right)+(k4\Leftrightarrow k_{5})+2\left(\frac{(k_{2}.k_{4})(k_{2}.k_{5})(k_{4}.k_{5})}{k_{2}^{2}\leavevmode\nobreak\ k_{4}^{2}\leavevmode\nobreak\ k_{5}^{2}}\right)\right]\\ &\times e^{i(k_{2}+k_{4}+k_{5}).r}\\ \end{split}$ (53)
$\propto\frac{G^{3}m_{1}m_{2}^{3}}{r^{11}}$
Integrals are done by starting with a simple integral ($I$) and taking its derivative with respect to $r$, where
$I$ $=$ $\int_{k}d^{3}k\frac{e^{ik.r}}{k^{2}}\propto\frac{1}{r}$ (54)
Similarly, we have other diagrams like
$\propto\frac{G^{3}m_{1}^{3}m_{2}}{r^{11}}$                         $\propto\frac{G^{3}m_{1}^{2}m_{2}^{2}}{r^{11}}$
$\propto\frac{G^{3}m_{1}^{2}m_{2}^{2}}{r^{11}}$
It was found that there are a total of 4 combinations for 2nd order $\phi^{3}$ diagrams. Only one case is considered in detail here, but all of them will either vanish or give the same fall of as $r^{-11}$.
$\sigma\phi^{2}$ $2^{nd}$ order
$\nabla^{2}\sigma_{ij}\partial_{i}\partial_{k}\phi\partial^{k}\partial^{j}\phi$ (55)
which in momentum space becomes $k_{1}^{2}\sigma(k_{1})_{ij}(k_{2}\cdot k_{3})k_{2}^{i}k_{3}^{j}$. Thus, we have shown that one of the interactions is very similar to the $2^{nd}$ order $\phi^{3}$ term. As a result, the calculations for the interactions involving the $2^{nd}$ order $\sigma\phi^{2}$ closely mirror the $2^{nd}$ order $\phi^{3}$ interactions, with the main difference being how the $\sigma$ field is incorporated into the diagrams. In $2^{nd}$ order $\sigma\phi^{2}$, the $\sigma$ field couples quadratically to $\phi$, which alters the diagram structure but does not affect the falloff of the interaction. The power law for distance dependence remains $\frac{1}{r^{11}}$, and the mass dependence in both cases is identical. Thus, while the couplings and diagrams may differ, the long-distance behavior and mass dependence follow the same pattern as in the $2^{nd}$ order $\phi^{3}$ theory. Hence, we have ignored the $2^{nd}$ order $\phi^{3}$ and $\sigma\phi^{2}$ contributions to the binding potential at 2 PN.